# Optimal Switching in Networked Control Systems: Finite Horizon

### 29 Mar 2026 v1

@@ Appendix E Proof of Theorem 4
$\mathbb{E}\!\Big[\big(S_{m}-C_{m}(x,\pi)\big)^{2}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Big]$
$\qquad\overset{(a)}{=}\mathbb{E}\!\Big[S_{m}^{2}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Big]+\big(C_{m}(x,\pi)\big)^{2}$
$\qquad\qquad-2\,C_{m}(x,\pi)\,\mathbb{E}\!\Big[S_{m}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Big]$
$\qquad\overset{(b)}{=}\sigma_{W}^{2}\sum_{j=0}^{m-1}a^{2(m-1-j)}+\big(C_{m}(x,\pi)\big)^{2}.$ (124)
In step (a), we used the fact that $C_{m}(x,\pi)$ is deterministic under the conditioning $X_{\ell-m}=x$. In step (b), we used the causality of the switching policy together with the independence and zero-mean property of $\{W_{k}\}$. Since $S_{m}=\sum_{j=0}^{m-1}a^{m-1-j}W_{\ell-m+j}$ depends only on the future disturbances $\{W_{\ell-m},\dots,W_{\ell-1}\}$, while $(X_{\ell-m},D_{\ell-m},Q_{\ell-m})$ is measurable with respect to $\sigma(W_{0},\dots,W_{\ell-m-1})$, it follows that
$\mathbb{E}\!\left[S_{m}\mid X_{\ell-m}=x,\;D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right]=0,$
@@
$\mathbb{E}\!\left[S_{m}^{2}\mid X_{\ell-m}=x,\;D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right]=\sigma_{W}^{2}\sum_{j=0}^{m-1}a^{2(m-1-j)}.$
@@
$\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\Bigg\{a^{2\tau}\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right)\mathbb{E}\!\Bigg[\big(S_{m}-C_{m}(X_{\ell-m},\pi)\big)^{2}\,\Big|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Bigg]\Bigg\}$
$=\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\Bigg\{a^{2\tau}\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right)\Bigg(\sigma_{W}^{2}\sum_{j=0}^{m-1}a^{2(m-1-j)}+\int_{-\infty}^{\infty}\big(C_{m}(x,\pi)\big)^{2}f^{\prime}_{X}(x)\,dx\Bigg)\Bigg\}$
$=\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\Bigg\{a^{2\tau}\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right)\int_{-\infty}^{\infty}\big(C_{m}(x,\pi)\big)^{2}f^{\prime}_{X}(x)\,dx\Bigg\},$ (125)
where the term $\sigma_{W}^{2}\sum_{j=0}^{m-1}a^{2(m-1-j)}$ is independent of the policy and therefore does not affect the minimization.
Since $\big(C_{m}(x,\pi)\big)^{2}\geq 0$ and $f^{\prime}_{X}(x)\geq 0$, the objective in (125) is minimized if and only if
@@ Appendix F Proof of Lemma 3
The third equality follows since, for $m\geq\tau$, the controller information $I^{\mathcal{C}}_{t_{k}+m+1}$ depends on switching decisions only up to $D_{t_{k}+m+1-\tau}$ and past disturbances, and is therefore independent of the current disturbance $W_{t_{k}+m}$. Hence, $\mathbb{E}[W_{t_{k}+m}\mid I^{\mathcal{C}}_{t_{k}+m+1}]=0$.
@@ Appendix G State-Based Policy
We derive the controller corresponding to the state-based switching policy

### 02 May 2026 v3

@@ Appendix E Proof of Theorem 4
$\mathbb{E}\!\Big[\big(S_{m}-C_{m}(X_{\ell-m},\pi)\big)^{2}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Big]$
$\qquad\overset{(a)}{=}\mathbb{E}\!\Big[S_{m}^{2}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Big]+\big(C_{m}(x,\pi)\big)^{2}$
$\qquad\qquad-2\,C_{m}(X_{\ell-m},\pi)\,\mathbb{E}\!\Big[S_{m}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Big]$
$\qquad\overset{(b)}{=}\mathbb{E}\!\Big[S_{m}^{2}+\big(C_{m}(X_{\ell-m},\pi)\big)^{2}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Big].$ (125)
In step (a), we used the fact that $C_{m}(x,\pi)$ is deterministic under the conditioning $X_{\ell-m}=x$. In step (b), we used the causality of the scheduling policy together with the independence and zero-mean property of $\{W_{k}\}$. Since $S_{m}=\sum_{j=0}^{m-1}a^{m-1-j}W_{\ell-m+j}$ depends only on the future disturbances $\{W_{\ell-m},\dots,W_{\ell-1}\}$, while $(X_{\ell-m},D_{\ell-m},Q_{\ell-m})$ is measurable with respect to $\sigma(W_{0},\dots,W_{\ell-m-1})$, it follows that
$\mathbb{E}\!\left[S_{m}C_{m}(X,\pi)\mid X_{\ell-m}=x,\;D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right]=0,$
@@
$\mathbb{E}\!\left[S_{m}^{2}\mid X_{\ell-m}=x,\;D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right]=\sigma_{W}^{2}\sum_{j=0}^{m-1}a^{2(m-1-j)}.$
@@
$\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\Bigg\{a^{2\tau}\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right)$
$\qquad\mathbb{E}\!\Bigg[\big(S_{m}-C_{m}(X_{\ell-m},\pi)\big)^{2}\,\Big|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Bigg]\Bigg\}$
$=\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\Bigg\{a^{2\tau}\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right)$
$\qquad\mathbb{E}\!\Bigg[S_{m}^{2}+C_{m}^{2}(X_{\ell-m},\pi)\,\Big|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Bigg]\Bigg\}$
$=\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}a^{2\tau}\mathbb{E}\!\Bigg[S_{m}^{2}+C_{m}^{2}(X_{\ell-m},\pi)\Bigg],$ (126)
Note that, for a fixed inter-update length $m$,
$\mathbb{E}[S_{m}^{2}\mid\Delta_{\ell}=m]$ $=\mathbb{E}\left[\left(\sum_{j=0}^{m-1}a^{m-1-j}W_{\ell-m+j}\right)^{2}\Bigg|\Delta_{\ell}=m\right]$
$=\sigma_{W}^{2}\sum_{j=0}^{m-1}a^{2(m-1-j)}$
$=\begin{cases}\dfrac{1-a^{2m}}{1-a^{2}}\sigma_{W}^{2},&\text{if }a\neq 1,\\[8.00003pt] m\sigma_{W}^{2},&\text{if }a=1.\end{cases}$ (127)
Therefore, after averaging over the random inter-update time $\Delta_{\ell}=\ell-t_{\ell}$, we obtain
$\mathbb{E}[S_{\Delta_{\ell}}^{2}]=\begin{cases}\dfrac{1-\mathbb{E}[a^{2\Delta_{\ell}}]}{1-a^{2}}\sigma_{W}^{2},&\text{if }a\neq 1,\\[8.00003pt] \mathbb{E}[\Delta_{\ell}]\sigma_{W}^{2},&\text{if }a=1.\end{cases}$ (128)
Hence, (126) becomes
$\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\;\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\right]=\begin{cases}\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\Bigg\{\dfrac{1-\mathbb{E}[a^{2\Delta_{\ell}}]}{1-a^{2}}\sigma_{W}^{2}+\mathbb{E}\!\left[C_{\Delta_{\ell}}^{2}(X_{t_{\ell}},\pi)\right]\Bigg\},&\text{if }a\neq 1,\\[10.00002pt] \min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\Bigg\{\mathbb{E}[\Delta_{\ell}]\sigma_{W}^{2}+\mathbb{E}\!\left[C_{\Delta_{\ell}}^{2}(X_{t_{\ell}},\pi)\right]\Bigg\},&\text{if }a=1.\end{cases}$ (129)
The first term depends only on the distribution of the inter-update time $\Delta_{\ell}$. In contrast, the second term depends on the conditional bias introduced by the policy through $C_{m}(X_{\ell-m},\pi).$ For any policy $\pi$ that induces a distribution $\Pr(\Delta_{\ell}=m)$ with $C_{m}(X_{\ell-m},\pi)\neq 0$, we can construct a randomized state-independent policy that induces the same distribution of $\Delta_{\ell}$. Since this randomized policy does not condition on the sign or magnitude of the state, it is symmetric and satisfies $C_{m}(X_{\ell-m},\pi)=0.$
Hence, since $\big(C_{m}(x,\pi)\big)^{2}\geq 0$ and $f^{\prime}_{X}(x)\geq 0$, the objective in (129) is minimized if and only if
@@ Appendix F Proof of Lemma 3
The third equality follows since, for $m\geq\tau$, the controller information $I^{\mathcal{C}}_{t_{k}+m+1}$ depends on scheduling decisions only up to $D_{t_{k}+m+1-\tau}$ and past disturbances, and is therefore independent of the current disturbance $W_{t_{k}+m}$. Hence, $\mathbb{E}[W_{t_{k}+m}\mid I^{\mathcal{C}}_{t_{k}+m+1}]=0$.
@@ Appendix G State-Based Policy
We derive the controller corresponding to the state-based scheduling policy
@@
Appendix H Extension to the Transient Case
In the transient case, the optimal controller in Thm. 1 becomes
$U_{k}=\gamma_{k}(I^{\mathcal{C}}_{k})=-L_{k}\,\hat{X}_{k},$ (143)
where
$\hat{X}_{k}=\mathbb{E}\bigl[X_{k}\mid I^{\mathcal{C}}_{k}\bigr],\quad L_{k}=\frac{ab\,P_{k+1}}{r+b^{2}P_{k+1}},\quad P_{k}=a^{2}P_{k+1}+q-\frac{(abP_{k+1})^{2}}{r+b^{2}P_{k+1}}.$ (144)
Hence, under this controller, the equivalent problem in Thm. 2 becomes
$P^{\prime}_{2}:\quad$ $\min_{\boldsymbol{\pi}\in\Pi}\;\frac{1}{N}\,\mathbb{E}\left[\sum_{k=0}^{N-1}L_{k}\bigl(r+b^{2}P_{k+1}\bigr)\mathcal{E}_{k}^{2}\right]$ (145)
s.t. $\frac{1}{N}\sum_{k=0}^{N-1}D_{k}\leq r_{s}.$
Compared to the steady-state case, the objective now includes an additional time-varying coefficient. This term propagates into the dynamic programming formulation and consequently appears in all subsequent equations.
First, the stage cost $g_{k}(I_{k}^{\mathcal{S}})$ in (33) is modified as follows:
$g_{k}(I_{k}^{\mathcal{S}})=\begin{cases}\displaystyle L_{k+\tau}(r+b^{2}P_{k+\tau+1})\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2},&\text{if }k=t_{k},\\[5.16663pt] \displaystyle L_{k+\tau}(r+b^{2}P_{k+\tau+1})\Bigg(a^{2\tau}\left(X_{k}-\mathbb{E}\!\left[X_{k}\mid I_{k+\tau}^{\mathcal{C}}\right]\right)^{2}+\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\Bigg),&\text{otherwise.}\end{cases}$ (146)
Accordingly, the DP recursion becomes
$V_{k}(I_{k}^{\mathcal{S}})=\min_{D_{k}}\Bigg\{$ $L_{k+\tau}(r+b^{2}P_{k+\tau+1})a^{2\tau}\bar{\mathcal{E}}_{k}^{2}+\mathbb{E}\!\left[V_{k+1}(I_{k+1}^{\mathcal{S}})\mid I_{k}^{\mathcal{S}},D_{k}=0\right],$
$\mathbb{E}\!\left[V_{k+1}(I_{k+1}^{\mathcal{S}})\mid I_{k}^{\mathcal{S}},D_{k}=1\right]\Bigg\}+L_{k+\tau}(r+b^{2}P_{k+\tau+1})\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}.$ (147)
The terminal-stage expressions in (38) and (39) become
$V^{\prime}_{(N-\tau-1)\,1\,1}(I_{N-\tau-1}^{\mathcal{S}})=L_{N-1}(r+b^{2}P_{N})\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2},$ (148)
and
$V^{\prime}_{(N-\tau-1)\,0\,0}(I_{N-\tau-1}^{\mathcal{S}})=L_{N-1}(r+b^{2}P_{N})\Big(a^{2\tau}\bar{\mathcal{E}}_{N-\tau-1}^{\,2}+\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\Big),$ (149)
respectively.
Hence, let $\ell=N-\tau-s$ for some $s\geq 1$. Then, Lemma 2 becomes
$V^{\prime}_{\ell\,s\,1}(I_{\ell}^{\mathcal{S}})=\left(\sum_{k=N-s}^{N-1}L_{k}(r+b^{2}P_{k+1})\right)\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right),$ (150)
and
$V^{\prime}_{\ell\,(s-1)\,0}(I_{\ell}^{\mathcal{S}})=L_{N-s}(r+b^{2}P_{N-s+1})a^{2\tau}\bar{\mathcal{E}}_{\ell}^{\,2}+\left(\sum_{k=N-s}^{N-1}L_{k}(r+b^{2}P_{k+1})\right)\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right),$ (151)
We now proceed to the separation proof. Since Lemma 2 has changed, the expression in (121) is modified as follows:
$=\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\;\Bigg\{\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right)$
$\qquad\qquad\times\mathbb{E}\!\Bigg[L_{N-s}(r+b^{2}P_{N-s+1})a^{2\tau}\left(S_{m}-C_{m}(X_{\ell-m},\pi)\right)^{2}\,\Bigg|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Bigg]$
$\qquad\qquad+\left(\sum_{k=N-s}^{N-1}L_{k}(r+b^{2}P_{k+1})\right)\bigg(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\bigg)\Bigg\}.$ (152)
Since the term
$\left(\sum_{k=N-s}^{N-1}L_{k}(r+b^{2}P_{k+1})\right)\bigg(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\bigg)$
is independent of $m$, and since $L_{N-s}(r+b^{2}P_{N-s+1})>0$, we recover the same expression as in (122). Therefore, symmetric policies remain optimal, and the separation principle continues to hold under the transient controller. We now proceed to the optimal scheduling mechanism.
With the modified DP recursion in (147), Lemma 4 becomes
where
$s_{(k-1)0}$ $\triangleq L_{k+\tau-1}(r+b^{2}P_{k+\tau})a^{2\tau}+a^{2}s_{k0},$ (154)
$c_{(k-1)00}$ $\triangleq s_{k0}+c_{k00}+L_{k+\tau-1}(r+b^{2}P_{k+\tau})\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}.$ (155)
Similarly, Lemma 5 becomes
where
$c_{(k-1)11}\triangleq s_{k0}+c_{k00}+L_{k+\tau-1}(r+b^{2}P_{k+\tau})\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\right).$ (157)
Finally, Lemmas 6 and 7 become
where
$c_{k(j-1)1}$ $\triangleq c_{(k+1)j0}\,\Pr\!\big(D_{k+1}=0\mid D_{k}=1\big)+c_{(k+1)j1}\,\Pr\!\big(D_{k+1}=1\mid D_{k}=1\big)$
$\qquad+L_{k+\tau}(r+b^{2}P_{k+\tau+1})\sum_{i=0}^{\tau-1}a^{2(\tau-1-i)}.$ (159)
and
where
$s_{(k-1)j}$ $\triangleq a^{2\tau}L_{k+\tau-1}(r+b^{2}P_{k+\tau})+a^{2}\Pr\!\big(D_{k}=0\mid D_{k-1}=0\big)s_{kj},$ (161)
$c_{(k-1)j0}$ $\triangleq\Pr\!\big(D_{k}=0\mid D_{k-1}=0\big)c_{kj0}+\Pr\!\big(D_{k}=1\mid D_{k-1}=0\big)c_{kj1}$
$\qquad+L_{k+\tau-1}(r+b^{2}P_{k+\tau})\sum_{i=0}^{\tau-1}a^{2(\tau-1-i)}.$ (162)
The terms $z_{k(j-1)1}$ and $z_{(k-1)j0}$ remain the same as in Lemmas 6 and 7, respectively.